OpenStudy (anonymous):

Not sure where to start for number 4 and 5. I think I got the first part of 4 right but unsure!! Any help is greatly appreciated!!!

OpenStudy (anonymous):

5. A 39,250 N force is applied to a 7200kg car initially at rest . what is the acceleration rate? How fast will the car be moving after 3,8 seconds? Sorry if you couldn't see the picture.

ganeshie8 (ganeshie8):

use newton's second law \(F=ma\)

OpenStudy (anonymous):

39,250 =7200 *a 39,250/7200 =a=5.5 m/s

ganeshie8 (ganeshie8):

careful about the units a=5.5 m/s^2

OpenStudy (anonymous):

ok. How do I calculate how fast it's going to be after 3.8 seconds. Do I use v=10 * t but that has something to od with free fall? So, no?

ganeshie8 (ganeshie8):

so the car is being accelerated for 3.8 seconds

ganeshie8 (ganeshie8):

\(a = 5.5\) \(t=3.8\) initial velocity, \(u=0\) final velocity, \(v=?\)

OpenStudy (anonymous):

5.58 * 3.8 =20.9 m/s

ganeshie8 (ganeshie8):

right, but it seems you have a rounding error.. maybe just enter below into the calculator : 39250/7200*3.8 = 20.7 m/s

OpenStudy (anonymous):

ok . I think I understand. Thank you so much!!!!

ganeshie8 (ganeshie8):

yw!

ganeshie8 (ganeshie8):

as you can see, "change in speed" requires acceleration no acceleration means no "change in speed"